# thanks: This letter has been submitted in conjunction with its companion manuscript Zhang et al. (2024)

### v1

@@ III Distribution-free goodness-of-fit
The residuals collected on the vector $\widehat{\bm{\epsilon}}$ allow us to quantify the discrepancy between the estimated angular power spectrum, $\widehat{\bm{C}}$, and the proposed model $\bm{C}^{M}(\bm{\theta})$. Therefore, to test the hypothesis in (4), we could rely solely on the process of partial sums
$w_{{}_{N}}(t)=\frac{1}{\sqrt{N}}\sum_{k\leq t}\widehat{\epsilon}_{k}=\frac{1}{\sqrt{N}}\widehat{\bm{\epsilon}}^{T}\mathbb{I}_{t}$ (5)
where $\mathbb{I}_{t}$ defines the $N\times 1$ vector of components $\mathbb{1}_{\{k\leq t\}}$, $\widehat{\epsilon}_{k}$, $k=1,\dots,N$ are the elements of $\widehat{\bm{\epsilon}}$, and each $k$ corresponds to a pair $(x_{i},\ell)$ – that is, $k=1$ indexes the pair $(x_{1},1)$, $k=L$ indexes $(x_{1},L)$, $k=L+1$ indexes $(x_{2},1)$, etc. A family of test statistics for (4) can then be easily constructed by considering functionals of $w_{{}_{N}}(t)$. For example, by choosing
$\begin{split}h_{\text{KS}}\bigl{[}w_{{}_{N}}(t)\bigl{]}&=\max_{t}|w_{{}_{N}}(t)|\end{split}$ (6)
@@
Regardless of the distribution of $\widehat{\bm{C}}$, under mild conditions (e.g., Philipp and Stout, 1975) and if $H_{0}$ holds, the process $w_{{}_{N}}(t)$ can be shown to converge to a Gaussian process with mean zero and covariance function depending on $\bm{C}^{M}(\bm{\theta})$ through the set of vectors $\{\bm{\mu}_{j}\}_{j=1}^{p}$ such that
@@
$\left[\begin{array}[]{c}\bm{\mu}^{T}_{1}\\ \vdots\\ \bm{\mu}^{T}_{p}\end{array}\right]=\left[\dot{\bm{C}}^{M}(\bm{\theta})^{T}\bm{\Sigma}^{-1}\dot{\bm{C}}^{M}(\bm{\theta})\right]^{-1/2}\left[\bm{\Sigma}^{-1/2}\dot{\bm{C}}^{M}(\bm{\theta})\right]^{T},$ (7)
where $p$ is the size of $\bm{\theta}$ and $\dot{\bm{C}}^{M}(\bm{\theta})=\frac{\partial\bm{C}^{M}(\bm{\theta})}{\partial\bm{\theta}}$ (see Zhang et al., 2024, for more details). We emphasize that the asymptotic normality of $w_{{}_{N}}(t)$ is a consequence of taking partial sums of weakly dependent quantities (Philipp and Stout, 1975), in our case, the residuals $\epsilon_{k}$. Hence, in our setup, large sample asymptotics means $N\rightarrow\infty$. The sample size can be further increased by considering multiple runs of the experiment or different time segments Zhang et al. (2024).
@@
The asymptotic null distribution of the statistics in (6) can be estimated with Monte Carlo or bootstrap methods. For example, let $\bm{C}^{M_{1}}(\bm{\theta})$ and $\bm{C}^{M_{2}}(\bm{\theta})$ be two candidate models for $E[\widehat{\bm{C}}]$ with components
$\begin{split}&C^{M_{1}}_{\ell}(x_{i},\bm{\theta})=\theta_{0}+\theta_{1}\ell+\theta_{2}x_{i},\\ &C^{M_{2}}_{\ell}(x_{i},\bm{\theta})=\exp\{\theta_{0}+\theta_{1}x_{i}+\theta_{2}x_{i}\ell\}.\end{split}$ (8)
@@
The $x_{i}$ values considered correspond to $n=100$ evenly-spaced points on the interval $[0,1]$, and $\ell=1,\dots,5$. The true value of the parameter $\bm{\theta}=(\theta_{0},\theta_{1},\theta_{2})$ is $(5,2,4)$. In the numerical experiments conducted, $\bm{\theta}$ is treated as unknown and is estimated as in (3). For each of the two models in (8), we consider two possible distributions for the corresponding estimator of the power spectrum:
@@
$\begin{split}&\widehat{\bm{C}}_{1,m}(x_{i})\sim\mathcal{N}\left(\bm{C}^{M_{m}}(x_{i},\bm{\theta}),\bm{\Sigma}(x_{i})\right),\\ &\widehat{\bm{C}}_{2,m}(x_{i})\sim T_{6}\left(\bm{C}^{M_{m}}(x_{i},\bm{\theta}),\bm{\Sigma}(x_{i})\right).\end{split}$ (9)
@@
That is, for each $x_{i}$, $\widehat{\bm{C}}_{1,m}(x_{i})$ and $\widehat{\bm{C}}_{2,m}(x_{i})$ are independent random vectors following, respectively, a multivariate normal and multivariate $t$-distribution with six degrees of freedom. The mean vectors, $\bm{C}^{M_{m}}(x_{i},\bm{\theta})$, have components $C^{M_{m}}_{\ell}(x_{i},\bm{\theta})$ defined as in (8) for $m=1,2$. The covariance matrices $\bm{\Sigma}(x_{i})$ are generated from a Wishart distribution with 10 degrees of freedom. The left panel of Figure 1 shows the simulated distribution function of the Kolmogorov-Smirnov statistic in (6) when the data is simulated from the four configurations described above and $H_{0}$ in (4) holds – that is, the model being tested and the one used to simulate the data are the same. Regardless of $\widehat{\bm{C}}_{1,m}(x_{i})$ and $\widehat{\bm{C}}_{2,m}(x_{i})$ being $t$ or Gaussian distributed, the null distribution of $h_{\text{KS}}[w_{{}_{N}}(t)]$ coincides when testing the same model. This demonstrates that whether or not $\widehat{\bm{C}}$ is Gaussian has no impact on the asymptotic null distribution of $h_{KS}[w_{{}_{N}}(t)]$. The model under study, however, does affect the asymptotic null distribution. As a result, one must simulate the null distribution of the desired test statistic on a case-by-case basis for each of the models being tested.
It is possible to overcome this limitation by constructing a new set of residuals, and a corresponding process of partial sums, with limiting null distribution independent of the model $\bm{C}^{M}(\bm{\theta})$ being tested. The main building block allowing us to perform this task is the operator
$U_{\bm{a},\bm{b}}\bm{q}=\bm{q}-\frac{\langle\bm{a}-\bm{b},\bm{q}\rangle}{1-\langle\bm{a},\bm{b}\rangle}(\bm{a}-\bm{b}),$
originally introduced in the context of distribution-free goodness-of-fit for discrete distributions Khmaladze (2013). Hereafter, we refer to the map entailed by the operator $U_{\bm{a},\bm{b}}$ as Khmaladze-2 (or K2) transformation  [^3] The nomenclature Khmaladze-2 (or K2) transformation is used to distinguish the map entailed by the operator $U_{\bm{a},\bm{b}}$ from the “Khmaladze transformation” or Khmaladze-1 (K1) transform proposed by the same author in Khmaladze (1982).
Given two vectors, $\bm{a}$ and $\bm{b}$, with $\|\bm{a}\|=\|\bm{b}\|=1$, the unitary operator $U_{\bm{a},\bm{b}}$ maps $\bm{q}=\bm{a}$ into $\bm{b}$, $\bm{q}=\bm{b}$ into $\bm{a}$, and leaves vectors $\bm{q}$ orthogonal to $\bm{a}$ and $\bm{b}$ unchanged. In our setup, $U_{\bm{a},\bm{b}}$ is employed to map the model-dependent vectors $\{\bm{\mu}_{j}\}_{j=1}^{p}$, characterizing the covariance of the process $w_{{}_{N}}(t)$, into an arbitrarily chosen, model-independent, orthonormal set of vectors $\{\bm{r}_{j}\}_{j=1}^{p}$. One such set can be obtained, for instance, by setting
$\bm{r}_{1}=\biggl{[}\frac{1}{\sqrt{N}},\dots,\frac{1}{\sqrt{N}}\biggl{]},$
@@
and constructing subsequent components via Gram-Schmidt orthonormalization of powers of a vector $\bm{r}_{2}$ with elements
$r_{2k}=\sqrt{\frac{12}{N}}\biggl{[}\frac{k}{N}-\frac{N+1}{2N}\biggl{]}.$
By adapting the testing framework of Khmaladze (2021) to our context, we can construct residuals with limiting null distribution depending solely on the vectors $\{\bm{r}_{j}\}_{j=1}^{p}$ and which is, consequently, independent of the model $\bm{C}^{M}(\bm{\theta})$ being tested. The main steps of this approach are now described.
  • Step 1. Estimate $\bm{\theta}$ via (3) and construct the vectors $\bm{\mu}_{j}$ in (7) with $\bm{\theta}$ replaced by $\widehat{\bm{\theta}}$.
  • Step 2. Given an arbitrarily chosen orthonormal set $\{\bm{r}_{j}\}_{j=1}^{p}$, construct a new set of vectors $\{\widetilde{\bm{r}}_{j}\}_{j=1}^{p}$ orthogonal to each $\bm{\mu}_{j^{\prime}}$, with $j^{\prime}<j$, by setting $\widetilde{\bm{r}}_{1}=\bm{r}_{1}$ and constructing subsequent vectors $\widetilde{\bm{r}}_{j}$ as
    $\widetilde{\bm{r}}_{j}=U_{\bm{\mu}_{(j-1)},\widetilde{\bm{r}}_{(j-1)}}\dots U_{\bm{\mu}_{1}\widetilde{\bm{r}}_{1}}\bm{r}_{j}$
    for example,
    $\widetilde{\bm{r}}_{2}$ $=U_{\bm{\mu}_{1}\widetilde{\bm{r}}_{1}}\bm{r}_{2}$
    $\widetilde{\bm{r}}_{3}$ $=U_{\bm{\mu}_{2}\widetilde{\bm{r}}_{2}}U_{\bm{\mu}_{1}\widetilde{\bm{r}}_{1}}\bm{r}_{3},\text{ etc.}$.
    Notice that each operator $U_{\bm{\mu}_{j},\widetilde{\bm{r}}_{j}}$ is applied to everything on its right-hand side.
  • Step 3. Consider the operator
    $\bm{U}_{p}^{T}\bm{q}=U_{\bm{\mu}_{1}\widetilde{\bm{r}}_{1}}\dots U_{\bm{\mu}_{p},\widetilde{\bm{r}}_{p}}\bm{q}$
    mapping vectors $\bm{\mu}_{j}$ into vectors $\bm{r}_{j}$ and construct the vector of K2 residuals $\widehat{\bm{e}}=\bm{U}_{p}^{T}\widehat{\bm{\epsilon}}$.
  • Step 4. Let the partial sums of the K2 residuals be
    $v_{{}_{N}}(t)=\frac{1}{\sqrt{N}}\widehat{\bm{e}}^{T}\mathbb{I}_{t}.$
    Similarly to (6), test statistics for (4) can be constructed by taking functionals $h[\cdot]$ of $v_{{}_{N}}(t)$.
  • Step 5. Compute the value of the desired test statistic, $h[v_{{}_{N}}(t)]$, and compare it with the distribution of the same functional taken with respect to a zero-mean Gaussian process, $u_{{}_{N}}(t)$, with covariance function $\langle u_{{}_{N}}(t),u_{{}_{N}}(t^{\prime})\rangle$ given by
    $\begin{split}\tfrac{1}{N}\mathbb{I}_{t}^{T}\biggl{[}I-\sum_{j=1}^{p}\bm{r}_{j}\bm{r}_{j}^{T}\biggl{]}\mathbb{I}_{t^{\prime}}.\\ \end{split}$ (10)
    The distribution of $h[u_{{}_{N}}(t)]$ can be derived via numerical simulation Zhang et al. (2024).
  • Step 6. Reject the model $\bm{C}^{M}(\bm{\theta})$ if the probability $P(h[u_{{}_{N}}(t)]>h[v_{{}_{N}}(t)])$ is smaller than the desired significance level.
The construction of the $\widetilde{\bm{r}}_{j}$ vectors in Step 2 is an intermediate step necessary to map vectors $\bm{\mu}_{j}$’s into $\bm{r}_{j}$ using the operator $\bm{U}_{p}^{T}$ Zhang et al. (2024). This enables the construction of the K2 transformed residuals, $\widehat{\bm{e}}$, as described in Step 3. In contrast to the process $w_{{}_{N}}(t)$ in (5), under $H_{0}$, the process of partial sums of the K2 residuals, $v_{{}_{N}}(t)$, constructed in Step 4, has the same limiting distribution as the Gaussian process $u_{{}_{N}}(t)$. The covariance function of the latter is given in (10) and depends solely on the -- arbitrarily chosen -- orthonormal set $\{\bm{r}_{j}\}_{j=1}^{p}$. This is true regardless of the distribution of $\widehat{\bm{C}}$ Zhang et al. (2024). It follows that, in the limit, test statistics constructed as outlined in Step 5 have a limiting null distribution that is independent of both the distribution of $\widehat{\bm{C}}$ and the model $\bm{C}^{M}(\bm{\theta})$ being tested. Finally, for sufficiently large $N$, the probability in Step 6 provides an asymptotic approximation of the p-value for testing the hypothesis in (4).
@@
To illustrate the validity of the proposed procedure with a numerical example, we consider the setup described in (8)-(9). The right panel of Figure 1 shows the simulated null distribution of the Kolmogorov-Smirnov statistic based on the K2 residuals,
$h_{\text{KS}}\bigl{[}v_{{}_{N}}(t)\bigl{]}=\max_{x}|v_{{}_{N}}(t)|,$ (11)
@@
for each of the four scenarios considered. All four curves effectively overlap with the simulated distribution of $h_{\text{KS}}\bigl{[}u_{{}_{N}}(t)]$ (blue-dashed lined). The latter represents the limit we expect to achieve, confirming the asymptotically distribution-free nature of the test.
Figure 2: A possible partition of a two-dimensional space within the scanning family. The black dots correspond to the points $\bm{x}_{i}$, the blue diamonds are the points $\bm{x}_{s}$ labelling the sets $[\bm{x}_{0},\bm{x}_{s}]$ and represented as grey squares. The point $\bm{x}_{0}$ is displayed as a red circle.
Finally, note that the approach described above easily extends to the case where $\bm{x}\in\mathcal{X}\subseteq\mathbb{R}^{D}$. Specifically, since each point $\bm{x}_{i}$ denotes a $D$-dimensional vector of components $(x_{1i},\dots,x_{Di})$, the partial sums are no longer taken over increasingly larger intervals but rather, can be taken over increasingly larger hyperrectangles
$[\bm{x}_{0},\bm{x}_{s}]=[x_{10},x_{1s}]\times\dots\times[x_{D0},x_{Ds}]$
with $x_{ds}\leq x_{ds^{\prime}}$, for $s\leq s^{\prime}$, for all $d=1,\dots,D$, and with $\bm{x}_{0}=(x_{10},\dots,x_{D0})$ denoting the point in $\mathcal{X}$ with coordinates given by the lower bounds for each of the $D$ dimensions considered.
One could construct a collection of sets satisfying this condition in many ways. For example, one could proceed by increasing one dimension at a time or increasing all dimensions simultaneously. Partitions of this type are known as a scanning family of sets (Khmaladze, 1993). One example is shown in Figure 2 where the black dots correspond to values $\bm{x}_{i}=(x_{i1},x_{i2})$, here taken equally spaced with respect to each dimension. The blue diamonds are the points $\bm{x}_{s}$, and the grey squares are the sets $[\bm{x}_{0},\bm{x}_{s}]$ labeled by such points. Hence, in this case, the “time” $t$ indexes the pairs $(\bm{x}_{s},\ell)$. That is, the partial sums of the residuals are taken over all the points falling within or on the edges of the grey squares, starting from the darkest square to the lightest, for each $\ell=1,\dots,L$.
IV Code availability

### v5

@@ III Distribution-free goodness-of-fit
The $N\times 1$ vector of decorrelated residuals, $\widehat{\mathbf{\epsilon}}$ in (8), allow us to quantify the discrepancy between the estimated angular power spectrum, $\widehat{\mathbf{C}}$, and the proposed model $\mathbf{C}^{M}(\mathbf{\theta})$. To test the hypothesis in (9), we consider the process of partial sums of the decorrelated residuals
$w_{{}_{N}}\hskip-0.71114pt(t)=\frac{1}{\sqrt{N}}\sum_{k=1}^{t}\widehat{\epsilon}_{k}=\frac{1}{\sqrt{N}}\widehat{\mathbf{\epsilon}}^{T}\mathbb{I}_{t}$ (10)
where $t=1,\dots,N$ indexes the elements of the process and $\mathbb{I}_{t}=[\mathbb{1}_{\{1\leq t\}},\mathbb{1}_{\{2\leq t\}},\dots,\mathbb{1}_{\{N\leq t\}}]^{T}$. Hence,
$w_{{}_{N}}(1)=\frac{\widehat{\epsilon}_{1}}{\sqrt{N}},w_{{}_{N}}(2)=\frac{\widehat{\epsilon}_{1}+\widehat{\epsilon}_{2}}{\sqrt{N}},\dots,w_{{}_{N}}(N)=\frac{1}{\sqrt{N}}\sum_{k=1}^{N}\widehat{\epsilon}_{k}.$ (11)
A family of test statistics for (9) is easily constructed by considering functionals of $w_{{}_{N}}\hskip-0.71114pt(t)$. For example, by choosing
$\begin{split}h_{\text{KS}}\bigl[w_{{}_{N}}\hskip-0.71114pt(t)\bigl]&=\max_{t}|w_{{}_{N}}\hskip-0.71114pt(t)|\end{split}$ (12)
@@
While the elements of $\widehat{\mathbf{\epsilon}}$ are uncorrelated, it is possible that some form of nonlinear dependence may persist among them when $\widehat{\mathbf{C}}$ is not approximately Gaussian. In most cases, however, this does not represent an impediment to deriving the limiting distribution of $w_{{}_{N}}\hskip-0.71114pt(t)$. In particular, if the dependence is ‘local’ as is the case, for example, when the vectors $\widehat{\mathbf{C}}^{T}(\mathbf{x}_{i})$, introduced in Section II, are independent from one another, under $H_{0}$ and as $N\rightarrow\infty$, the process $w_{{}_{N}}\hskip-0.71114pt(t)$ can be shown to converge to a Gaussian process with mean zero and covariance function approximately equal to:
$\frac{1}{N}\mathbb{I}^{T}_{t}[\mathbf{I}_{N}-\sum_{j=1}^{p}\mathbf{\mu}_{j}\mathbf{\mu}^{T}_{j}\bigl]\mathbb{I}_{t^{\prime}}$ (13)
with $\mathbf{I}_{{}_{N}}$ denoting the $N\times N$ identity matrix and the set of vectors $\{\mathbf{\mu}_{j}\}_{j=1}^{p}$ is such that:
@@
$\left[\begin{array}[]{c}\mathbf{\mu}^{T}_{1}\\ \vdots\\ \mathbf{\mu}^{T}_{p}\end{array}\right]=\left[\dot{\mathbf{C}}^{M}(\mathbf{\theta})^{T}\bm{\Sigma}^{-1}\dot{\mathbf{C}}^{M}(\mathbf{\theta})\right]^{-1/2}\left[\bm{\Sigma}^{-1/2}\dot{\mathbf{C}}^{M}(\mathbf{\theta})\right]^{T},$ (14)
where $p$ is the size of $\mathbf{\theta}$ and $\dot{\mathbf{C}}^{M}(\mathbf{\theta})=\frac{\partial\mathbf{C}^{M}(\mathbf{\theta})}{\partial\mathbf{\theta}}$. The structure of the vectors $\mathbf{\mu}_{j}$ strictly depends on the type of estimator employed (see Zhang et al. (2025) for more details).
From (13), it follows that the asymptotic null distribution of $w_{{}_{N}}\hskip-0.71114pt(t)$ depends on the model being tested through its covariance. As a result, one must simulate the null distribution of the desired test statistic on a case-by-case basis for each of the models being tested with Monte Carlo methods or the parametric bootstrap [^2] The parametric bootstrap is a resampling method that differs from classical Monte Carlo methods in that various samples are simulated by replacing the true value of $\mathbf{\theta}$ with its estimator $\widehat{\mathbf{\theta}}$ Tibshirani and Efron (1993).. Unfortunately, this may be computationally expensive when testing a number of different, and possibly complicated, models. Nevertheless, it is possible to overcome this limitation by constructing a new set of residuals and a corresponding process of partial sums, with limiting null distribution independent of $\mathbf{C}^{M}(\mathbf{\theta})$.
The main building block is the operator
$U_{\mathbf{a},\mathbf{b}}\mathbf{q}=\mathbf{q}-\frac{\langle\mathbf{a}-\mathbf{b},\mathbf{q}\rangle}{1-\langle\mathbf{a},\mathbf{b}\rangle}(\mathbf{a}-\mathbf{b}),$ (15)
originally introduced in the context of distribution-free goodness-of-fit for discrete distributions Khmaladze (2013). Hereafter, we refer to the map entailed by the operator $U_{\mathbf{a},\mathbf{b}}$ as Khmaladze-2 (or K2) transformation  [^3] The nomenclature Khmaladze-2 (or K2) transformation is used to distinguish the map entailed by the operator $U_{\mathbf{a},\mathbf{b}}$ from the “Khmaladze transformation” or Khmaladze-1 (K1) transformation Khmaladze (1982)..
Given two vectors, $\mathbf{a}$ and $\mathbf{b}$, with $\|\mathbf{a}\|=\|\mathbf{b}\|=1$, the unitary operator $U_{\mathbf{a},\mathbf{b}}$ maps $\mathbf{q}=\mathbf{a}$ into $\mathbf{b}$, $\mathbf{q}=\mathbf{b}$ into $\mathbf{a}$, and leaves vectors $\mathbf{q}$ orthogonal to $\mathbf{a}$ and $\mathbf{b}$ unchanged. In our setup, $U_{\mathbf{a},\mathbf{b}}$ is employed to map the model-dependent vectors $\{\mathbf{\mu}_{j}\}_{j=1}^{p}$, characterizing the covariance of the process $w_{{}_{N}}\hskip-0.71114pt(t)$, into an arbitrarily chosen, model-independent, orthonormal set of vectors $\{\mathbf{r}_{j}\}_{j=1}^{p}$.
By adapting the testing framework of Khmaladze (2021) to our context, we can construct residuals with limiting null distribution depending solely on the vectors $\{\mathbf{r}_{j}\}_{j=1}^{p}$ and which is, consequently, independent of the model $\mathbf{C}^{M}(\mathbf{\theta})$ being tested. The main steps of this approach are now described.
  • Step 1. Estimate $\mathbf{\theta}$ via (7) and construct the vectors $\mathbf{\mu}_{j}$ in (14) with $\mathbf{\theta}$ replaced by $\widehat{\mathbf{\theta}}$.
  • Step 2. Given an arbitrarily chosen orthonormal set $\{\mathbf{r}_{j}\}_{j=1}^{p}$, construct a new set of vectors $\{\widetilde{\mathbf{r}}_{j}\}_{j=1}^{p}$ orthogonal to each $\mathbf{\mu}_{j^{\prime}}$, with $j^{\prime}<j$, by setting $\widetilde{\mathbf{r}}_{1}=\mathbf{r}_{1}$ and constructing subsequent vectors $\widetilde{\mathbf{r}}_{j}$ as
    $\widetilde{\mathbf{r}}_{j}=U_{\mathbf{\mu}_{(j-1)},\widetilde{\mathbf{r}}_{(j-1)}}\dots U_{\mathbf{\mu}_{1}\widetilde{\mathbf{r}}_{1}}\mathbf{r}_{j}$ (16)
    for example,
    $\begin{split}\widetilde{\mathbf{r}}_{2}&=U_{\mathbf{\mu}_{1}\widetilde{\mathbf{r}}_{1}}\mathbf{r}_{2}\\ \widetilde{\mathbf{r}}_{3}&=U_{\mathbf{\mu}_{2}\widetilde{\mathbf{r}}_{2}}U_{\mathbf{\mu}_{1}\widetilde{\mathbf{r}}_{1}}\mathbf{r}_{3},\text{ etc.}\end{split}$ (17)
    Notice that each operator $U_{\mathbf{\mu}_{j},\widetilde{\mathbf{r}}_{j}}$ is applied to everything on its right-hand side.
  • Step 3. Consider the operator
    $\mathbf{U}_{p}^{T}\mathbf{q}=U_{\mathbf{\mu}_{1}\widetilde{\mathbf{r}}_{1}}\dots U_{\mathbf{\mu}_{p},\widetilde{\mathbf{r}}_{p}}\mathbf{q}$ (18)
    mapping vectors $\mathbf{\mu}_{j}$ into vectors $\mathbf{r}_{j}$ and construct the vector of K2 residuals
    $\widehat{\mathbf{e}}=\mathbf{U}_{p}^{T}\Bigl[\widehat{\mathbf{\epsilon}}-\sum_{j=1}^{p}\mathbf{\mu}_{j}\mathbf{\mu}^{\sf T}_{j}\widehat{\mathbf{\epsilon}}\Bigl]=\mathbf{U}_{p}^{T}\widehat{\mathbf{\epsilon}}-\sum_{j=1}^{p}\mathbf{r}_{j}\mathbf{r}^{\sf T}_{j}\mathbf{U}_{p}^{T}\widehat{\mathbf{\epsilon}}$ (19)
  • Step 4. Let the partial sums of the K2 residuals be
    $v_{{}_{N}}\hskip-0.71114pt(t)=\frac{1}{\sqrt{N}}\widehat{\mathbf{e}}^{T}\mathbb{I}_{t}.$ (20)
    Similarly to (12), test statistics for (9) can be constructed by taking functionals $h[\cdot]$ of $v_{{}_{N}}\hskip-0.71114pt(t)$.
  • Step 5. Compute the value of the desired test statistic, $h[v_{{}_{N}}\hskip-0.71114pt(t)]$, and compare it with the distribution of the same functional taken with respect to a zero-mean Gaussian process, $u_{{}_{N}}\hskip-0.71114pt(t)$, with covariance function
[…]
    The distribution of $h[u_{{}_{N}}\hskip-0.71114pt(t)]$ can be derived directly by simulating realizations of
    $u_{{}_{N}}\hskip-0.71114pt(t)=\frac{1}{\sqrt{N}}\mathbb{I}_{t}^{T}[\mathbf{e}-\sum_{j=1}^{p}\mathbf{r}_{j}\mathbf{r}^{T}_{j}\mathbf{e}]$ (22)
    with $\mathbf{e}$ drawn from a multivariate standard normal.
  • Step 6. Reject the model $\mathbf{C}^{M}(\mathbf{\theta})$ if the probability $P(h[u_{{}_{N}}\hskip-0.71114pt(t)]>h[v_{{}_{N}}\hskip-0.71114pt(t)])$ is smaller than the desired significance level.
The construction of the $\widetilde{\mathbf{r}}_{j}$ vectors in Step 2 is an intermediate step necessary to map vectors $\mathbf{\mu}_{j}$’s into $\mathbf{r}_{j}$ using the operator $\mathbf{U}_{p}^{T}$ Zhang et al. (2025). This enables the construction of the K2 transformed residuals, $\widehat{\mathbf{e}}$, as described in Step 3. In contrast to the process $w_{{}_{N}}\hskip-0.71114pt(t)$ in (10), under $H_{0}$, the process of partial sums of the K2 residuals, $v_{{}_{N}}\hskip-0.71114pt(t)$, constructed in Step 4, has the same limiting distribution as the Gaussian process $u_{{}_{N}}\hskip-0.71114pt(t)$. The covariance function of the latter is given in (21) and depends solely on the - arbitrarily chosen - orthonormal set $\{\mathbf{r}_{j}\}_{j=1}^{p}$. This is true regardless of the distribution of $\widehat{\mathbf{C}}$ Zhang et al. (2025). It follows that, in the limit, test statistics constructed as outlined in Step 5 have a limiting null distribution that is independent of both the distribution of $\widehat{\mathbf{C}}$ and the model $\mathbf{C}^{M}(\mathbf{\theta})$ being tested. Finally, for sufficiently large $N$, the probability in Step 6 provides an asymptotic approximation of the p-value for testing the hypothesis in (9).
@@
IV An illustrative example
Let $\mathbf{C}^{M_{1}}(\mathbf{\theta})$ and $\mathbf{C}^{M_{2}}(\mathbf{\theta})$ be two candidate models for $E[\widehat{\mathbf{C}}$] with components
$\begin{split}&C^{M_{1}}_{\ell}(\mathbf{x}_{i},\mathbf{\theta})=\theta_{0}+\theta_{1}\ell+\theta_{2}\mathbf{x}_{i},\\ &C^{M_{2}}_{\ell}(\mathbf{x}_{i},\mathbf{\theta})=\exp\{\theta_{0}+\theta_{1}\mathbf{x}_{i}+\theta_{2}\mathbf{x}_{i}\ell\}.\end{split}$ (23)
@@ IV An illustrative example
The $\mathbf{x}_{i}$ values considered correspond to $n=100$ evenly-spaced points on the interval $[0,1]$, and $\ell=1,\dots,5$. The true value of the parameter $\mathbf{\theta}=(\theta_{0},\theta_{1},\theta_{2})$ is $(5,2,4)$. In the numerical experiments conducted, $\mathbf{\theta}$ is treated as unknown and is estimated as in (7). For each of the two models in (23), we consider two possible distributions for the corresponding estimator of the power spectrum:
@@
$\begin{split}&\widehat{\mathbf{C}}_{m,1}(\mathbf{x}_{i})\sim\mathcal{N}\left(\mathbf{C}^{M_{m}}(\mathbf{x}_{i},\mathbf{\theta}),\bm{\Sigma}(\mathbf{x}_{i})\right),\\ &\widehat{\mathbf{C}}_{m,2}(\mathbf{x}_{i})\sim T_{6}\left(\mathbf{C}^{M_{m}}(\mathbf{x}_{i},\mathbf{\theta}),\bm{\Sigma}(\mathbf{x}_{i})\right).\end{split}$ (24)
@@
That is, for each $\mathbf{x}_{i}$, $\widehat{\mathbf{C}}_{m,1}(\mathbf{x}_{i})$ and $\widehat{\mathbf{C}}_{m,2}(\mathbf{x}_{i})$ are independent random vectors following, respectively, a multivariate normal and multivariate $t$-distribution with six degrees of freedom. The mean vectors, $\mathbf{C}^{M_{m}}(\mathbf{x}_{i},\mathbf{\theta})$, have components $C^{M_{m}}_{\ell}(\mathbf{x}_{i},\mathbf{\theta})$ defined as in (23) for $m=1,2$. The covariance matrices $\bm{\Sigma}(\mathbf{x}_{i})$ are generated from a Wishart distribution with 10 degrees of freedom. The left panel of Figure 1 shows the simulated distribution function of the Kolmogorov-Smirnov statistic in (12) when the data is simulated from the four configurations described above and $H_{0}$ in (9) holds – that is, the model being tested and the one used to simulate the data are the same. Regardless of $\widehat{\mathbf{C}}_{m,1}(\mathbf{x}_{i})$ and $\widehat{\mathbf{C}}_{m,2}(\mathbf{x}_{i})$ being $t$ or Gaussian distributed, the null distribution of $h_{\text{KS}}[w_{{}_{N}}\hskip-0.71114pt(t)]$ coincides when testing the same model. This demonstrates that whether or not $\widehat{\mathbf{C}}$ is Gaussian has no impact on the asymptotic null distribution of $h_{KS}[w_{{}_{N}}\hskip-0.71114pt(t)]$. As expected, however, the model under study does affect the asymptotic null distribution.
Let us now proceed to implement the procedure described in Steps 1-6. The orthonormal set $\{\mathbf{r}_{j}\}_{j=1}^{p}$ selected is such that
$\mathbf{r}_{1}=\biggl[\frac{1}{\sqrt{N}},\dots,\frac{1}{\sqrt{N}}\biggl],$ (25)
@@
and subsequent components are constructed via Gram-Schmidt orthonormalization of powers of a vector $\mathbf{r}_{2}$ with elements
$r_{2k}=\sqrt{\frac{12}{N}}\biggl[\frac{k}{N}-\frac{N+1}{2N}\biggl].$ (26)
@@
The right panel of Figure 1 shows the simulated null distribution of the Kolmogorov-Smirnov statistic based on the K2 residuals,
$h_{\text{KS}}\bigl[v_{{}_{N}}\hskip-0.71114pt(t)\bigl]=\max_{t}|v_{{}_{N}}\hskip-0.71114pt(t)|,$ (27)
@@
for each of the four scenarios considered. All four curves effectively overlap with the simulated distribution of $h_{\text{KS}}\bigl[u_{{}_{N}}\hskip-0.71114pt(t)]$ (blue-dashed lined). The latter represents the limit we expect to achieve, confirming the asymptotically distribution-free nature of the test.
In summary, these results suggest that one can simulate the null distribution of $h_{\text{KS}}[w_{{}_{N}}\hskip-0.71114pt(t)]$ assuming that $\widehat{\mathbf{C}}$ is Gaussian, even when it is not, without affecting the result of the test. Such a distribution, however, will differ when testing different models, and thus would require a separate simulation for each of them. Alternatively, one could proceed with a statistical test based on $h_{\text{KS}}\bigl[u_{{}_{N}}\hskip-0.71114pt(t)]$. In this case, the null distribution would be the same, regardless of the model being tested, and could be easily obtained numerically as summarized in Step 5.
V Code availability